# Exact electromagnetic multipole expansion using elementary current multipoles

### v1

@@ Abstract
Multipole expansion plays an important role in the description of electromagnetic scatterers, allowing them to be accurately characterized by a small set of expansion coefficients. However, to describe electromagnetic excitations inside a scatterer, the current density in it should be decomposed into current multipoles. Such current expansion includes nonradiating current configurations that are absent in the classical field-based expansion. Unfortunately, the use of current multipoles has so far been limited by the absence of an exact and general expression for the current multipole moments beyond their point-multipole approximation. Here, for the first time to our knowledge, we derive such an expression and present the exact mapping relations between the classical and current multipole moments. We use our theory to calculate the scattering and extinction cross sections for large, wavelength-scale, optical scatterers supporting multipole excitations up to the sixth order. Our results show perfect agreement with the Mie theory. The expressions are valid for electromagnetic scatterers of arbitrary sizes and shapes without restrictions on the multipole orders. They describe all possible current configurations, including those that do not produce any fields in the far-field zone. Current multipoles complement the existing theory for electromagnetic multipole expansion, while their minimalistic and universal character makes them a convenient tool for characterizing and designing diverse electromagnetic scattering systems of arbitrary complexity.
@@ I Introduction
The radiation and scattering of electromagnetic fields are among the most fundamental phenomena in physics. Their rigorous theoretical treatment is essential in many areas of science and technology, from telecommunications [1] to atmospheric science [2, 3] and astrophysics [4]. Recently, scattering theory has become particularly relevant for the rapidly growing field of nanophotonics, in which the resonant optical response of nanostructured media has created new paradigms for optical field manipulation and enhanced light-matter interactions [5, 6]. For example, Mie multipole resonances [7, 8, 9, 10, 11] have been extensively explored for the realization of various optical phenomena, including Kerker effect [12, 13, 14], toroidal and anapole excitations [15, 16], and bound states in the continuum [17, 18, 19]. The engineering of multipole excitations has also been used in various functional structures, such as antireflection coatings [20], Huygens metasurfaces [21, 22], nanoscale lasers [23], metamaterials [24, 25], as well as systems exhibiting structural colors [26], collective resonances [27, 28, 29, 30], and enhanced nonlinear optical effects [31, 32].
Describing and designing resonant micro- and nanoscatterers can be done using electromagnetic multipole expansion. It allows one to characterize every finite-sized scattering system with a relatively small set of expansion coefficients corresponding to mutually orthogonal electromagnetic fields of different-order multipoles [33]. Since the original Mie solutions [34], multipole expansion has been approached in many ways [35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. However, regardless of the approach, the classical field-based expansion does not fully capture the actual current oscillations within the system. For example, electromagnetic anapoles are absent in this expansion, as they do not radiate into the far field [45, 46, 47, 48], so that their theoretical description requires the artificial separation of the toroidal parts of the multipole moments [49, 43, 11]. Moreover, if the scatterer is a part of a larger scattering system, its own scattered field is difficult to isolate. Therefore, the current oscillations within that scatterer cannot be easily determined through the classical field-based expansion [37, 50]. Finally, the multipole moments in the Cartesian field-based expansion [40] lack a generalized expression for an arbitrary order $l$, which limits their use to structures supporting only low-order multipole modes.
To overcome these limitations, a complementary multipole expansion has been proposed [37]. It expands the scattering current density
$\textbf{J}(\textbf{r})=-i\omega\epsilon_{0}[\epsilon_{\text{r}}(\mathbf{r})-\epsilon_{\text{s}}]\textbf{E}(\textbf{r})$ (1)
@@
similarly to the moments of the primitive Cartesian multipoles [51]. In the above equations, $\epsilon_{\text{r}}(\mathbf{r})$ is the position-dependent relative electric permittivity and $\epsilon_{\text{s}}$ is its value in the surrounding medium, $l$ is the multipole order, and $\mathbf{J}(\mathbf{r})\textbf{r}^{l-1}$ is the outer product of vector $\mathbf{J}(\mathbf{r})$ and l1l-1 vectors $\mathbf{r}$. The current multipoles are not divided into electric and magnetic multipoles and represent very simple configurations of linear current elements (see Fig. 1), which facilitates their use in designing individual and arrayed scatterers with prescribed scattering characteristics [52, 53, 48, 50, 54, 55].
@@
Figure 1: Examples of elementary current configurations in the excitations of a current dipole ($l$ = 1), quadrupole ($l$ = 2), octupole ($l$ = 3), hexadecapole ($l$ = 4), and triacontadipole ($l$ = 5), with moments $p_{x}$, $Q_{yx}$, $O_{zxy}$, $H_{xxyz}$, and $T_{xxxyz}$, respectively. Current multipoles of high orders, such as 4 and 5, are relatively common, since often, they are constituents of classical multipoles of lower orders. In particular, current hexadecapoles are constituents of classical electric quadrupoles and magnetic octupoles, whereas current triacontadipoles appear in classical electric octupoles and magnetic hexadecapoles; see Eqs. (24)-(42). In numerical calculations presented in this work, we consider classical multipoles up to the electric and magnetic hexacontatetrapoles ($l$ = 6) [56] which include current multipoles of orders $l$ up to 8.
Unfortunately, Eq. (2) is not valid beyond the long-wavelength (point-multipole) approximation and cannot be used to evaluate the multipole moments and their contributions to the scattering and extinction cross sections for wavelength-scale and larger scatterers. So far, the only method for obtaining the exact current multipole moments for such scatterers was to calculate them from the exact moments of the classical electric and magnetic multipoles, i.e., from the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$, where $l$ and $m$ are the expansion orders [37, 50, 54, 55]. However, because current multipoles include nonradiating current configurations that are absent in the classical multipole expansion, the mapping from the classical to current multipole moments is underdetermined, so that some of the current multipole moments cannot be unambiguously evaluated. The lack of a complete theoretical framework for the current multipole expansion has so far prevented researchers from using the approach to its full potential, despite its clear advantages over the classical multipole expansion.
where $(2l-1)!!=(2l-1)\cdot(2l-3)\cdot...\cdot 5\cdot 3\cdot 1$ is the double factorial, $j_{l-1}(kr)$ are the spherical Bessel functions of the first kind, and $k$ is the wavenumber in the surrounding medium. The derivation of this expression is presented in Appendix A. Equation (3) allows the moments, i.e., the elements of the tensor $\bm{M}^{(l)}_{\text{exact}}$, to be calculated directly and unambiguously from the scattering current density distribution $\mathbf{J}(\mathbf{r})$ defined in Eq. (1). Several examples of these multipole moments are presented in Table 1, and the corresponding configurations of linear current elements in the point-multipole approximation are shown schematically in Fig. 1. Expressions in Table 1 illustrate the fact that every current multipole moment corresponds to a unique monomial containing a Cartesian vector component of the scattering current density ($J_{x}$, $J_{y}$, $J_{z}$) and, for $l\geq 2$, a product of normalized Cartesian coordinates ($x/r$, $y/r$, $z/r$). These moments can be called elementary as they constitute the simplest Cartesian configurations of oscillating currents. At the same time, they are exact and applicable to scatterers of arbitrary sizes and shapes, which we demonstrate in the next section.
@@
Table 1: Exact expressions for the current multipole moments illustrated schematically in Fig. 1. The names of the multipoles follow the same convention as in Ref. [57].
@@
II Numerical verification
To verify the correctness of Eq. (3), we used it to numerically calculate the multipole contributions to the scattering and absorption cross sections for wavelength-scale scatterers across the visible and near-infrared spectral range. We did this by calculating the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ in terms of the elements of tensor $\bm{M}^{(l)}_{\text{exact}}$ and then evaluated their contributions to the cross sections in question.
@@ II Numerical verification
where $C_{1}=-ik^{3}/(6\pi\epsilon E_{0})$, $C_{2}=-k^{4}/(60\pi\epsilon E_{0})$, and $C_{3}=-ik^{5}/(210\pi\epsilon E_{0})$, with $E_{0}$ being the incident field amplitude and $\epsilon=\epsilon_{0}\epsilon_{\text{s}}$. The relations for a set of higher-order multipoles are presented in Appendix B. Additionally, in Ref. [58] we provide a Python code to automatically derive expressions for an arbitrary order $l$.
@@
$Q_{\text{ext,E}}(l)=-\frac{\pi}{k^{2}S}\sum_{m=-1,+1}(2l+1)\operatorname{Re}[ma_{\text{E}}(l,m)],$ (9)
$Q_{\text{ext,M}}(l)=-\frac{\pi}{k^{2}S}\sum_{m=-1,+1}(2l+1)\operatorname{Re}[a_{\text{M}}(l,m)],$ (10)
where $S$ is the geometric cross section of the scatterer ($S=\pi R^{2}$ for a spherical nanoparticle of radius $R$). The corresponding multipole contributions to the normalized absorption cross sections are then calculated as
@@
In the numerical calculations, $\mathbf{J}(\mathbf{r})$ was obtained using Eq. (1) from the electric field distribution calculated in Comsol Multiphysics for a silicon sphere of diameter 600 nm (Fig. 2) and a silver sphere of diameter 400 nm (Fig. 3). In both cases, we assumed the scatterers to be embedded in PMMA, which is a dielectric material commonly used in nanofabrication. The optical constants of Si, Ag, and PMMA were taken from Refs. [59, 60, 61, 62]. The scatterers were illuminated by a plane wave linearly polarized along the $x$ axis and propagating along the $z$ axis, which is appropriate for the extinction cross sections defined by Eqs. (9) and (10). Figures 2 and 3 show the calculated cross sections. The points correspond to the approach based on the numerical evaluation of the current multipole moments using Eq. (3), while the lines show the direct predictions of the Mie theory. The approaches are seen to be in excellent agreement.

### v2

@@ Abstract
Multipole expansion plays an important role in the description of electromagnetic scatterers, allowing them to be accurately characterized by a small set of expansion coefficients. However, to describe electromagnetic excitations inside a scatterer, the current density in it should be decomposed into current multipoles, which include nonradiating current configurations (anapoles) that are absent in the classical field-based expansion. Unfortunately, the use of current multipoles has so far been limited by the absence of an exact and general expression for the current multipole moments beyond their point-multipole approximation. Here, we derive such an expression and present the exact mapping relations between the classical and current multipole moments. We use our theory to calculate the scattering and extinction cross sections for large, wavelength-scale, optical scatterers supporting multipole excitations up to the sixth order, showing perfect agreement with the Mie theory. We also demonstrate the ability of current multipole expansion to describe anapole excitations beyond the small-scatterer approximation, which allows us to derive the exact anapole condition and reveal the actual current configurations and their contributions to scattering. Our theoretical framework is valid for electromagnetic scatterers of arbitrary sizes and shapes without restrictions on the multipole orders, complementing the existing theory of electromagnetic multipole expansion. The minimalistic and universal character of current multipoles makes them a convenient tool for characterizing and designing diverse electromagnetic scattering systems of arbitrary complexity.
@@ I Introduction
The radiation and scattering of electromagnetic fields are among the most fundamental phenomena in physics. Their rigorous theoretical treatment is essential in many areas of science and technology, from telecommunications [5] to atmospheric science [52, 8] and astrophysics [27]. Recently, scattering theory has become particularly relevant for the rapidly growing field of nanophotonics, in which the resonant optical response of nanostructured media has created new paradigms for optical field manipulation and enhanced light-matter interactions [31, 56]. For example, multipole excitations at Mie resonances [30, 38, 37, 41, 3] have been extensively explored for the realization of Kerker effect [10, 43, 60], toroidal and anapole excitations [42, 66, 11, 9], bound states in the continuum [35, 53, 12], antireflection coatings [64], Huygens metasurfaces [13, 40], nanoscale lasers [65], metamaterials [50, 22], structural colors [4], nonlocal (collective) resonances [18, 39, 33, 32], and enhanced nonlinear optical effects [63, 54].
Theoretical description and design of the scattering properties of micro- and nanoscale objects can be done using electromagnetic multipole expansion. It allows one to characterize every finite-sized scattering system with a relatively small set of expansion coefficients corresponding to mutually orthogonal electromagnetic fields of different-order multipoles [28]. Since the original Mie theory [45], multipole expansion has been approached in many ways [47, 16, 24, 17, 15, 1, 2, 14, 25, 44, 48, 19]. However, regardless of the approach, the classical field-based expansion does not fully capture the actual current oscillations within the system. For example, electromagnetic anapoles are absent in this expansion, as they do not radiate into the far field [55, 6, 36, 62]. As a result, theoretical description of anapoles often involves the so-called toroidal moments, which are artificially separated parts of the actual multipole moments [46, 25, 3]. Since the conventional multipole expansion is based on the radiated fields rather than their sources (i.e., oscillating currents), it is difficult to use it for engineering the electromagnetic scatterers.
To overcome these limitations, a complementary multipole expansion has been proposed [24]. It expands the scattering current density $\textbf{J}(\textbf{r})$ in the scatterer, defined as
$\textbf{J}(\textbf{r})=-i\omega\epsilon_{0}[\epsilon_{\text{r}}(\mathbf{r})-\epsilon_{\text{s}}]\textbf{E}(\textbf{r}),$ (1)
@@
similarly to the moments of the primitive Cartesian multipoles [26]. In the above equations, $\epsilon_{\text{r}}(\mathbf{r})$ is the position-dependent relative electric permittivity and $\epsilon_{\text{s}}$ is its value in the surrounding medium, $l$ is the multipole order, and $\mathbf{J}(\mathbf{r})\textbf{r}^{l-1}$ is the outer product of vector $\mathbf{J}(\mathbf{r})$ and l1l-1 vectors $\mathbf{r}$. The current multipoles are not divided into electric and magnetic multipoles and represent very simple configurations of linear current elements (see Fig. 1), which facilitates their use in designing individual and arrayed scatterers with prescribed near-field characteristics [23, 61, 62, 57, 58, 59].
@@
Figure 1: Examples of current configurations in the excitations of current multipoles of different orders $l$ in the point-multipole approximation: a current dipole ($l$ = 1), quadrupole ($l$ = 2), octupole ($l$ = 3), hexadecapole ($l$ = 4), and triacontadipole ($l$ = 5), corresponding to moments $p_{x}$, $Q_{yx}$, $O_{zxy}$, $H_{xxyz}$, and $T_{xxxyz}$, respectively. Each arrow represents a single linear current element (see Ref. [24]). Current multipoles of high orders, such as 4 and 5, are relatively common, constituting classical multipoles of lower orders. In particular, current hexadecapoles are constituents of classical electric quadrupoles and magnetic octupoles, whereas current triacontadipoles appear in classical electric octupoles and magnetic hexadecapoles (see Eqs. (47)-(65)). In numerical calculations presented in this work, we consider classical multipoles up to the electric and magnetic hexacontatetrapoles ($l$ = 6) which include current multipoles of orders $l$ up to 8. Note that, in the used convention, the outermost current element at positive coordinates is positive if the multipole moment is positive.
@@
Table 1: Exact expressions for the current multipole moments illustrated schematically in Fig. 1. The names of the multipoles follow the same convention as in Refs. [21, 49].
@@
Unfortunately, Eq. (2) is not valid beyond the long-wavelength (point-multipole) approximation and cannot be used to accurately evaluate the magnitudes of the multipole moments and their contributions to the scattering and extinction cross sections for wavelength-scale and larger scatterers. So far, the only method for obtaining the exact current multipole moments for such scatterers was to calculate them from the exact moments of the classical electric and magnetic multipoles, i.e., from the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$, where $l$ and $m$ are the expansion orders [24, 22, 57, 58, 59]. However, because current multipoles include nonradiating current configurations that are absent in the classical multipole expansion, the mapping from the classical to current multipole moments is underdetermined, so that some of the current multipole moments cannot be unambiguously evaluated [57]. The lack of a complete theoretical framework for the current multipole expansion has so far prevented researchers from using the approach to its full potential, despite its clear advantages that make it complementary to the classical multipole expansion.
Figure 2: Numerical validation of the exact expression for the current multipole moments [Eq. (3)] using the optical response of a silicon sphere of diameter 600 nm embedded in PMMA, for which we evaluate the multipole contributions to the normalized scattering cross section $Q_{\text{scat}}$ (a, b) and absorption cross section $Q_{\text{abs}}=Q_{\text{ext}}-Q_{\text{scat}}$ (c, d) defined in Eqs. (8)-(11). The dots correspond to the contributions of moments $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$ obtained by numerically calculating the current multipole moments $M^{(l)}_{\text{exact}}$ in Comsol Multiphysics using Eqs. (1) and (3) and converting them to the classical multipole moments using Eqs. (4)-(7), (47)-(65), and Ref. [34]. The solid lines are the cross sections obtained from the Mie theory. The insets show the peaks of electric and magnetic triacontadipoles ($l=5$) and hexacontatetrapoles ($l=6$). In the main plots, the curves for subsequent multipoles are vertically offset for clarity, and some of them are magnified by a constant factor, as indicated. The results of numerical multipole expansion and Mie theory overlap perfectly (up to the numerical precision of Comsol), which validates the expressions presented in this work.
where $(2l-1)!!=(2l-1)\cdot(2l-3)\cdot...\cdot 5\cdot 3\cdot 1$ is the double factorial, $j_{l-1}(kr)$ are the spherical Bessel functions of the first kind, and $k$ is the wavenumber in the surrounding medium. The derivation of this expression is presented in Appendix A. Equation (3) allows the moments, i.e., the elements of the tensor $\bm{M}^{(l)}_{\text{exact}}$, to be calculated directly and unambiguously from the scattering current density distribution $\mathbf{J}(\mathbf{r})$ defined in Eq. (1). Moreover, the moments can be used to assess the magnitude of the corresponding current density (see Appendix B and Refs. [58, 59]). Several examples of these moments are presented in Table 1, and the corresponding configurations of linear current elements in the point-multipole approximation are shown schematically in Fig. 1. The expressions in Table 1 illustrate the fact that every current multipole moment corresponds to a unique monomial containing a Cartesian vector component of the scattering current density ($J_{x}$, $J_{y}$, $J_{z}$) and, for $l\geq 2$, a product of normalized Cartesian coordinates $x/r$, $y/r$, and $z/r$. These moments constitute exceptionally simple Cartesian configurations of oscillating currents, which makes them truly the elementary multipole moments. At the same time, they are exact and applicable to scatterers of arbitrary sizes and shapes, which we demonstrate in the next section.
Figure 3: Numerical validation of the current multipole expansion using the optical response of a silver sphere of diameter 400 nm embedded in PMMA. Similarly to Fig. 2, the solid lines represent the cross sections calculated using the Mie theory, while the dots correspond to the cross sections obtained numerically in Comsol Multiphysics using the current multipole expansion (Eqs. (1), (3), (4)-(7), (47)-(65), and Ref. [34]). In the given example, the scattering cross section ($Q_{\text{scat}}$) has significant contributions of (a) the electric multipoles of order $l\leq 6$ (up to hexacontatetrapoles) and (b) the magnetic multipoles of order $l\leq 3$ (up to octupoles), while the absorption cross section ($Q_{\text{abs}}$) is dominated by the electric multipoles of $l\geq$4, as shown in (c). Similarly to Fig. 2, we find an excellent agreement between the numerical multipole expansion and Mie theory.
II Numerical validation
To validate Eq. (3), we used it to numerically calculate the multipole contributions to the scattering and absorption cross sections for wavelength-scale scatterers across the visible and near-infrared spectral range. We did this by first calculating the elements of tensor $\bm{M}^{(l)}_{\text{exact}}$, and then using them to calculate the multipole expansion coefficients $a_{\text{E}}(l,m)$ and $a_{\text{M}}(l,m)$. Finally, we evaluated their contributions to the cross sections in question.
@@ II Numerical validation
where $C_{1}=-ik^{3}/(6\pi\epsilon E_{0})$, $C_{2}=-k^{4}/(60\pi\epsilon E_{0})$, and $C_{3}=-ik^{5}/(210\pi\epsilon E_{0})$, with $E_{0}$ being the incident field amplitude and $\epsilon=\epsilon_{0}\epsilon_{\text{s}}$. The mapping relations for a set of higher-order multipoles are presented in Appendix C. Additionally, in Ref. [34] we provide a Python code to automatically derive the mapping relations for an arbitrary order $l$.
@@
[…]
where $S$ is the geometric cross section of the scatterer ($S=\pi R^{2}$ for a spherical nanoparticle of radius $R$). The corresponding multipole contributions to the normalized absorption cross section are then calculated as
@@
In the numerical calculations, $\mathbf{J}(\mathbf{r})$ defined in Eq. (1) was obtained from the electric field distribution calculated in Comsol Multiphysics for a silicon sphere of diameter 600 nm (see Fig. 2) and a silver sphere of diameter 400 nm (see Fig. 3). In both cases, we assumed the scatterers to be embedded in PMMA, which is a dielectric material commonly used in nanofabrication. The optical constants of Si, Ag, and PMMA were taken from Refs. [51, 20, 29, 7]. The scatterers were illuminated by a plane wave linearly polarized along the $x$ axis and propagating along the $z$ axis, which corresponds to the definition of the extinction cross section given in Eqs. (9) and (10). In the calculation of multipole moments, the coordinate system had its origin at the center of the scatterer. Figures 2 and 3 show the calculated cross sections. In the presented plots, the dots were obtained using the current multipole moments computed numerically in Comsol Multiphysics with the aid of Eq. (3) together with Eqs. (1), (4)-(11), (47)-(65) and Ref. [34], while the solid lines were calculated directly using the Mie theory. The results obtained with both approaches are seen to be in excellent agreement.
In the considered examples, the scatterers are comparable in size to the wavelength and cannot be treated as point scatterers. Using the multipole expansion in the long-wavelength (small-scatterer) approximation for such scatterers is known to produce significant errors [1]. Hence, the presented analysis of wavelength-scale scatterers confirms the validity of our theory beyond the long-wavelength approximation, demonstrating its applicability to scatterers of an arbitrary size.